# Shenzhen, China

### v3

@@ 3.3 Image Generation performance
Fig. 6b shows the generation effect of the sample image. Fig. 6c shows the retrieval accuracy using generated images. The generated images have high semantic similarity with the seen images and have good diversity in low-level visual features, which can be manipulated by the guidance scale hyperparameter (Fig. 6d). Surprisingly, the retrieval accuracy using generated images is even higher than that using EEG embeddings, implying the benefit of our two-stage visual reconstruction for image retrieval (See Appendix 0.C for Figure 11). We also report the decoding and reconstruction performance for EEG, MEG, and fMRI across various metrics in the Appendix 0.D.
@@ 0.C.2 Stage II - CLIP-Embedded Image Synthesis
In Figure 11, we compare the effects of one-stage and two-stage EEG-guided image generation. We show images generated using EEG embeddings directly (One-stage) and images generated using image embeddings obtained via prior diffusion (Two-stage). It can be seen that the two-stage EEG-guided image generation can more accurately reconstruct the semantic and low-level visual features of the original image, and the style is more realistic.
@@ 0.G.1 Accuracy for time windows
EEGITNet
83.23 $\pm$ 2.77
65.67 $\pm$ 3.84
43.87 $\pm$ 3.78
7.79 $\pm$ 1.27
22.12 $\pm$ 2.88
EEGConformer
89.18 $\pm$ 2.88
75.49 $\pm$ 5.30
55.32 $\pm$ 6.88
11.29 $\pm$ 3.55
33.49 $\pm$ 6.78
ShallowFBCSPNet
84.29 $\pm$ 2.56
67.13 $\pm$ 3.61
46.07 $\pm$ 4.55
8.26 $\pm$ 2.95
25.87 $\pm$ 5.03
EEGNetV4
89.03 $\pm$ 2.86
75.40 $\pm$ 5.18
56.77 $\pm$ 6.60
13.29 $\pm$ 3.99
35.50 $\pm$ 7.14
B.D.
91.14 $\pm$ 2.45
79.59 $\pm$ 4.59
62.62 $\pm$ 6.33
16.29 $\pm$ 4.35
42.16 $\pm$ 7.56
NICE
92.17 $\pm$ 2.78
82.10 $\pm$ 5.55
65.65 $\pm$ 8.03
19.40 $\pm$ 5.83
46.26 $\pm$ 10.42
MLP
87.67 $\pm$ 4.41
72.58 $\pm$ 7.62
53.52 $\pm$ 9.45
11.81 $\pm$ 4.33
32.74 $\pm$ 9.24
ATM-S (Ours)
93.89 $\pm$ 1.97
85.38 $\pm$ 3.95
71.63 $\pm$ 6.17
22.84 $\pm$ 5.87
52.22 $\pm$ 8.34
@@
EEGITNet
77.14 $\pm$ 4.14
55.29 $\pm$ 6.22
31.88 $\pm$ 6.20
2.93 $\pm$ 1.70
13.00 $\pm$ 4.45
EEGConformer
79.49 $\pm$ 3.91
59.31 $\pm$ 5.83
36.67 $\pm$ 5.35
4.11 $\pm$ 1.83
17.04 $\pm$ 4.21
ShallowFBCSPNet
75.14 $\pm$ 4.45
53.06 $\pm$ 6.70
31.23 $\pm$ 6.56
3.19 $\pm$ 2.17
13.01 $\pm$ 5.15
EEGNetV4
82.60 $\pm$ 3.17
64.28 $\pm$ 5.44
42.24 $\pm$ 6.10
6.13 $\pm$ 2.40
21.23 $\pm$ 5.19
B.D.
81.49 $\pm$ 3.52
62.35 $\pm$ 6.44
40.53 $\pm$ 7.19
6.16 $\pm$ 2.40
20.45 $\pm$ 5.40
NICE
81.85 $\pm$ 2.53
63.57 $\pm$ 4.46
41.86 $\pm$ 4.68
6.43 $\pm$ 1.46
21.39 $\pm$ 3.62
MLP
80.49 $\pm$ 2.61
62.09 $\pm$ 3.88
40.37 $\pm$ 3.71
5.67 $\pm$ 1.09
19.97 $\pm$ 3.02
ATM-S (Ours)
82.88 $\pm$ 5.18
65.03 $\pm$ 8.51
44.86 $\pm$ 9.85 8.04 $\pm$ 3.34 24.62 $\pm$ 8.17
ATM-E (Ours) 83.30 $\pm$ 3.92 65.80 $\pm$ 6.98
44.84 $\pm$ 7.76
7.47 $\pm$ 2.84
23.75 $\pm$ 6.86
@@
EEGITNet
76.69 $\pm$ 12.97
56.98 $\pm$ 16.31
36.35 $\pm$ 15.11
5.75 $\pm$ 3.62
18.14 $\pm$ 9.40
EEGConformer
76.17 $\pm$ 13.13
56.29 $\pm$ 16.70
34.72 $\pm$ 14.79
3.98 $\pm$ 2.80
17.10 $\pm$ 9.21
ShallowFBCSPNet
74.32 $\pm$ 12.14
53.97 $\pm$ 15.81
33.48 $\pm$ 14.35
6.10 $\pm$ 4.61
16.53 $\pm$ 9.94
EEGNetV4
92.81 $\pm$ 2.22
83.15 $\pm$ 4.20
67.81 $\pm$ 6.11
19.51 $\pm$ 5.19
48.99 $\pm$ 6.75
B.D.
78.42 $\pm$ 8.81
58.24 $\pm$ 12.13
37.97 $\pm$ 11.38
5.88 $\pm$ 3.49
18.61 $\pm$ 7.81
NICE
92.73 $\pm$ 2.75
83.26 $\pm$ 5.47
67.96 $\pm$ 8.31
19.32 $\pm$ 5.33
49.26 $\pm$ 9.69
MLP
83.09 $\pm$ 2.54
66.70 $\pm$ 4.16
45.43 $\pm$ 4.58
7.23 $\pm$ 1.66
25.14 $\pm$ 3.66
@@
EEGITNet
77.47 $\pm$ 3.75
55.75 $\pm$ 6.01
33.60 $\pm$ 6.16
3.80 $\pm$ 2.01
14.02 $\pm$ 3.89
EEGConformer
68.02 $\pm$ 7.75
44.24 $\pm$ 9.25
23.84 $\pm$ 7.65
1.66 $\pm$ 1.37
8.62 $\pm$ 4.20
ShallowFBCSPNet
76.40 $\pm$ 4.64
54.81 $\pm$ 7.09
32.37 $\pm$ 7.28
2.49 $\pm$ 1.57
13.24 $\pm$ 5.98
EEGNetV4
82.60 $\pm$ 3.17
64.28 $\pm$ 5.44
42.24 $\pm$ 6.10
6.13 $\pm$ 2.40
21.23 $\pm$ 5.19
B.D.
85.30 $\pm$ 6.24
69.87 $\pm$ 10.09
50.82 $\pm$ 11.74
10.57 $\pm$ 5.26
29.74 $\pm$ 11.28
NICE
83.75 $\pm$ 3.21
65.82 $\pm$ 5.93
44.41 $\pm$ 6.12
7.04 $\pm$ 2.83
23.55 $\pm$ 4.86
MLP
80.49 $\pm$ 2.61
62.09 $\pm$ 3.88
40.37 $\pm$ 3.71
5.67 $\pm$ 1.09
19.97 $\pm$ 3.02
ATM-S (Ours)
87.36 $\pm$ 3.97
72.80 $\pm$ 7.02
53.80 $\pm$ 8.41 11.84 $\pm$ 4.80 33.73 $\pm$ 8.73
ATM-E (Ours) 87.41 $\pm$ 3.07 72.85 $\pm$ 5.51
53.15 $\pm$ 6.52
11.12 $\pm$ 3.26
32.61 $\pm$ 6.78

### v4

@@ 3.3 Image Generation performance
Fig. 6b shows the generation effect of the sample image. Fig. 6c shows the retrieval accuracy using generated images. The generated images have high semantic similarity with the seen images and have good diversity in low-level visual features, which can be manipulated by the guidance scale hyperparameter (Fig. 6d). Surprisingly, the retrieval accuracy using generated images is even higher than that using EEG embeddings, implying the benefit of our two-stage visual reconstruction for image retrieval (See Appendix 0.C for Fig.11). We also report the decoding and reconstruction performance for EEG, MEG, and fMRI across various metrics in the Appendix 0.D.
@@ 0.C.2 Stage II - CLIP-Embedded Image Synthesis
In Fig. 11, we compare the effects of one-stage and two-stage EEG-guided image generation. We show images generated using EEG embeddings directly (One-stage) and images generated using image embeddings obtained via prior diffusion (Two-stage). It can be seen that the two-stage EEG-guided image generation can more accurately reconstruct the semantic and low-level visual features of the original image, and the style is more realistic.
@@ 0.G.1 Accuracy for time windows
EEGITNet 83.23 $\pm$ 2.77 65.67 $\pm$ 3.84 43.87 $\pm$ 3.78 7.79 $\pm$ 1.27 22.12 $\pm$ 2.88
EEGConformer 89.18 $\pm$ 2.88 75.49 $\pm$ 5.30 55.32 $\pm$ 6.88 11.29 $\pm$ 3.55 33.49 $\pm$ 6.78
ShallowFBCSPNet 84.29 $\pm$ 2.56 67.13 $\pm$ 3.61 46.07 $\pm$ 4.55 8.26 $\pm$ 2.95 25.87 $\pm$ 5.03
EEGNetV4 89.03 $\pm$ 2.86 75.40 $\pm$ 5.18 56.77 $\pm$ 6.60 13.29 $\pm$ 3.99 35.50 $\pm$ 7.14
B.D. 91.14 $\pm$ 2.45 79.59 $\pm$ 4.59 62.62 $\pm$ 6.33 16.29 $\pm$ 4.35 42.16 $\pm$ 7.56
NICE 92.17 $\pm$ 2.78 82.10 $\pm$ 5.55 65.65 $\pm$ 8.03 19.40 $\pm$ 5.83 46.26 $\pm$ 10.42
MLP 87.67 $\pm$ 4.41 72.58 $\pm$ 7.62 53.52 $\pm$ 9.45 11.81 $\pm$ 4.33 32.74 $\pm$ 9.24
ATM-S (Ours) 93.89 $\pm$ 1.97 85.38 $\pm$ 3.95 71.63 $\pm$ 6.17 22.84 $\pm$ 5.87 52.22 $\pm$ 8.34
@@
EEGITNet 77.14 $\pm$ 4.14 55.29 $\pm$ 6.22 31.88 $\pm$ 6.20 2.93 $\pm$ 1.70 13.00 $\pm$ 4.45
EEGConformer 79.49 $\pm$ 3.91 59.31 $\pm$ 5.83 36.67 $\pm$ 5.35 4.11 $\pm$ 1.83 17.04 $\pm$ 4.21
ShallowFBCSPNet 75.14 $\pm$ 4.45 53.06 $\pm$ 6.70 31.23 $\pm$ 6.56 3.19 $\pm$ 2.17 13.01 $\pm$ 5.15
EEGNetV4 82.60 $\pm$ 3.17 64.28 $\pm$ 5.44 42.24 $\pm$ 6.10 6.13 $\pm$ 2.40 21.23 $\pm$ 5.19
B.D. 81.49 $\pm$ 3.52 62.35 $\pm$ 6.44 40.53 $\pm$ 7.19 6.16 $\pm$ 2.40 20.45 $\pm$ 5.40
NICE 81.85 $\pm$ 2.53 63.57 $\pm$ 4.46 41.86 $\pm$ 4.68 6.43 $\pm$ 1.46 21.39 $\pm$ 3.62
MLP 80.49 $\pm$ 2.61 62.09 $\pm$ 3.88 40.37 $\pm$ 3.71 5.67 $\pm$ 1.09 19.97 $\pm$ 3.02
ATM-S (Ours) 82.88 $\pm$ 5.18 65.03 $\pm$ 8.51 44.86 $\pm$ 9.85 8.04 $\pm$ 3.34 24.62 $\pm$ 8.17
ATM-E (Ours) 83.30 $\pm$ 3.92 65.80 $\pm$ 6.98 44.84 $\pm$ 7.76 7.47 $\pm$ 2.84 23.75 $\pm$ 6.86
@@
EEGITNet 76.69 $\pm$ 12.97 56.98 $\pm$ 16.31 36.35 $\pm$ 15.11 5.75 $\pm$ 3.62 18.14 $\pm$ 9.40
EEGConformer 76.17 $\pm$ 13.13 56.29 $\pm$ 16.70 34.72 $\pm$ 14.79 3.98 $\pm$ 2.80 17.10 $\pm$ 9.21
ShallowFBCSPNet 74.32 $\pm$ 12.14 53.97 $\pm$ 15.81 33.48 $\pm$ 14.35 6.10 $\pm$ 4.61 16.53 $\pm$ 9.94
EEGNetV4 92.81 $\pm$ 2.22 83.15 $\pm$ 4.20 67.81 $\pm$ 6.11 19.51 $\pm$ 5.19 48.99 $\pm$ 6.75
B.D. 78.42 $\pm$ 8.81 58.24 $\pm$ 12.13 37.97 $\pm$ 11.38 5.88 $\pm$ 3.49 18.61 $\pm$ 7.81
NICE 92.73 $\pm$ 2.75 83.26 $\pm$ 5.47 67.96 $\pm$ 8.31 19.32 $\pm$ 5.33 49.26 $\pm$ 9.69
MLP 83.09 $\pm$ 2.54 66.70 $\pm$ 4.16 45.43 $\pm$ 4.58 7.23 $\pm$ 1.66 25.14 $\pm$ 3.66
@@
EEGITNet 77.47 $\pm$ 3.75 55.75 $\pm$ 6.01 33.60 $\pm$ 6.16 3.80 $\pm$ 2.01 14.02 $\pm$ 3.89
EEGConformer 68.02 $\pm$ 7.75 44.24 $\pm$ 9.25 23.84 $\pm$ 7.65 1.66 $\pm$ 1.37 8.62 $\pm$ 4.20
ShallowFBCSPNet 76.40 $\pm$ 4.64 54.81 $\pm$ 7.09 32.37 $\pm$ 7.28 2.49 $\pm$ 1.57 13.24 $\pm$ 5.98
EEGNetV4 82.60 $\pm$ 3.17 64.28 $\pm$ 5.44 42.24 $\pm$ 6.10 6.13 $\pm$ 2.40 21.23 $\pm$ 5.19
B.D. 85.30 $\pm$ 6.24 69.87 $\pm$ 10.09 50.82 $\pm$ 11.74 10.57 $\pm$ 5.26 29.74 $\pm$ 11.28
NICE 83.75 $\pm$ 3.21 65.82 $\pm$ 5.93 44.41 $\pm$ 6.12 7.04 $\pm$ 2.83 23.55 $\pm$ 4.86
MLP 80.49 $\pm$ 2.61 62.09 $\pm$ 3.88 40.37 $\pm$ 3.71 5.67 $\pm$ 1.09 19.97 $\pm$ 3.02
ATM-S (Ours) 87.36 $\pm$ 3.97 72.80 $\pm$ 7.02 53.80 $\pm$ 8.41 11.84 $\pm$ 4.80 33.73 $\pm$ 8.73
ATM-E (Ours) 87.41 $\pm$ 3.07 72.85 $\pm$ 5.51 53.15 $\pm$ 6.52 11.12 $\pm$ 3.26 32.61 $\pm$ 6.78